OpenStudy (redbirdd):

Solve for the roots in the following equation. Hint: Factor both quadratic expressions. (x^4 + 5x^2 - 36)(2x^2 + 9x - 5) = 0

OpenStudy (3mar):

May I help?

OpenStudy (otherworldly):

go ahead 3mar

OpenStudy (mww):

The two factors to begin with each can be factorised. First one, the quartic, best to convert into a quadratic form Second one, factorise as a quadratic trinomial.

OpenStudy (redbirdd):

I have no idea what you're talking about. I haven't even been taught this. I basically just get it thrown at me and get crappy "help" from my teacher.

OpenStudy (3mar):

Sorry for late.

OpenStudy (3mar):

Let's begin with left parentheses.

OpenStudy (3mar):

Can you factorize this: \[x^4+5x^2-36=0\]?

OpenStudy (mww):

I don't think they know. For that one best to let u = x^2 and convert it to a quadratic then x^4 + 5x^2 - 36 becomes u^2 + 5u - 36

OpenStudy (3mar):

@redbirdd, Are you there?

OpenStudy (3mar):

where is she to proceed with you!?

OpenStudy (mww):

the method itself is not exactly straight forward. Here's a link that can help. The method is known as the 'product, sum factor' (PSF) method. https://www.mathsisfun.com/algebra/factoring-quadratics.html

OpenStudy (3mar):

It is the rule and it is applicable for all such problems. \[(x^4+5x^2-36)(2x^2+9x-5)=0\] \[(x^2+9)(x^2-4)(x+5)(2x-1)=0\] \[(x^2+9)(x-2)(x+2)(x+5)(2x-1)=0\] each parentheses is equal to zero, then you can find the roots::: \[x^2+9=0\] \[x^2=-9\] \[x=\pm \sqrt{-9}=\pm (\sqrt{-1}*\sqrt{9})=\pm 3i\] where \[i=\sqrt{-1}\] is an imaginary number \[x-2=0\] \[x=2\] \[x+2=0\] \[x=-2\] \[x+5=0\] \[x=-5\] \[2x-1=0\] \[x=\frac{ 1 }{ 2 }\] So the solution is \[{3 , -3 , -5 , \frac{ 1 }{ 2 } }\] if x belongs to real numbers and it is \[{3 , -3 , -5 , \frac{ 1 }{ 2 } , 3i , -3i}\] for x belongs to imaginary numbers

OpenStudy (3mar):

Thank you for the medal! mmw